4 workers can build a small barn in 20 days. How many workers are needed if we want to build 5 in the same length of time (20 days)? Submit Answer Tries 0/10 How many workers are needed if we want to accelerate construction and build one barn in 4 days? Submit Answer Tries 0/10 On the other hand, if we want to build 5 barns in 4 days, how many workers do we need?
Answer #1

Question 1:

Answer = (5 * 4) = 20

Question 2:

Answer = (4 * 20 / 4) = 20

Question 3:

Answer = (5 * 4) * (4 * 20 / 4) = 400

** if the answer does not match please comment